You invested $12,000 in two accounts paying 6% and 8% annual interest, respectively. If the total interest earned for the year was $880, how much was invested at each rate?
Found 2 solutions by mananth, richwmiller:
Answer by mananth(16946)

Part I 8.00% per annum ------------- Amount invested =x
Part II 6.00% per annum ------------ Amount invested = y
12000
Interest----- 880.00

Part I 8.00% per annum ---x
Part II 6.00% per annum ---y
Total investment
x + 1 y= 12000 -------------1
Interest on both investments
8.00% x + 6.00% y= 880
Multiply by 100
8 x + 6 y= 88000.00 --------2
Multiply (1) by -8
we get
-8 x -8 y= -96000.00
Add this to (2)
0 x -2 y= -8000
divide by -2
y = 4000
Part I 8.00% $ 8000
Part II 6.00% $ 4000

CHECK
8000 --------- 8.00% ------- 640.00
4000 ------------- 6.00% ------- 240.00
Total -------------------- 880.00

m.ananth@hotmail.ca

Answer by richwmiller(17219)

x+y=12000,
0.06*x+0.08*y=880
x=12000-y
0.06*(12000-y)+0.08*y=880
720-0.06y+0.08*y=880
0.02*y=160
y=8000 at 8%
x=12000-y
x=4000 at 6%
check
0.06*4000+0.08*8000=880
240+640=880
880=880
ok